# Information conservation in de Sitter tunneling

### v1

@@ 2.1 Painlevé coordinate and time-like geodesic line equation
By substituting (52) and (53) into (13), we obtain the Painlevé coordinate:
@@
Obviously, in (54), the coordinate singularity at the event horizon is eliminated. The selection of positive and negative signs in (54) will have a clearer physical meaning in the equation of time-like geodesics.
In this article, we discuss particles that carry both electric and magnetic charges. So they are all massive particles, not following the optical geodesic equation. We consider massive particles as de Broglie s-waves, employing an approach analogous to that described in ref. 27 . The time-like geodesic equation, derived from the WKB approximation, is presented below:
@@
Where $k$ is de Broglie wave number. By substituting $g_{00}$ and $g_{01}$ in (54) into (55), we can obtain:
@@
It can be clearly seen that when taking a positive sign, (58) is the equation of motion for the outgoing particles near the outer event horizon of the black hole, while when taking a negative sign, (58) is the equation of motion for the incoming particles near the cosmic horizon. Furthermore, if considering the influence of self gravity, $M$ in (54) and (58) should be replaced by $M\mp\omega$, and $Q_{h}$ in (54) and (58) should be replaced by $Q_{h}\mp q_{h}$. $\omega$ is the energy of the particle, $q_{h}$ is the equivalent charge of the particle. The minus sign and plus sign represent the outgoing particles and incoming particles, respectively.
@@ 2.2 Tunneling out of the black hole outer event horizon
According to (60), the imaginary part of the action can be written as:
@@
To proceed with our calculations, we substitute the Hamilton’s equation into (62):
@@
By substituting (63) and (64) into (62), we can obtain:
@@
Where $E_{Q_{h}}$ is the energy corresponding to the electromagnetic field. $M$, $E_{Q_{h}}$ is a fixed and unchanging quantity. $M-\omega$, $E_{Q_{h}-q_{h}}$ is the residual gravitational and electromagnetic energy of the black hole after radiation. It can be clearly seen that in our formula, energy conservation, charge conservation, and magnetic charge conservation are all reflected. It is worth emphasizing that in order to simplify the equivalent charge of a black hole after radiation to $Q_{h}-q_{h}$, we adopt the following assumption:
@@
Where $q_{e}$ and $q_{g}$ are the charges and magnetic charges of the emitted particles, respectively. The loss of gravitational energy and electromagnetic energy caused by particle tunneling in a black hole is represented by the following equations:
@@
By substituting (67) and (68) into (65), we can obtain:
@@
Substituting (70) into (69), we have:
@@ 2.3 Tunneling into the cosmic horizon
(78) and (79) respectively represent the changes in gravitational energy and electromagnetic energy after particles enter the cosmic horizon, from which it can be clearly seen that energy conservation, charge conservation, and magnetic charge conservation occur. Substituting (76), (78) and (79) into (77) yields:
@@ 2.4 Discussion
Comparing (86) and (87) with (72) and (81) respectively, the temperature $T^{\prime}$ at the event horizon surface of the black hole and the temperature $T^{\prime\prime}$ at the cosmic horizon surface are obtained:
@@ 3 Kerr-Newman-Kasuya de Sitter black hole
Let $A=\frac{3}{\Lambda}(a^{2}+Q_{h}^{2})$, $B=\frac{3}{\Lambda}-a^{2}$. The equation (96) has four solutions, $r_{non}$, $r_{-}$, $r_{+}$, $r_{c}$. Among them, $r_{non}$ is a negative value solution with no physical meaning, $r_{-}$ is the inner event horizon of black hole, $r_{+}$ is the outer event horizon of black hole, and $r_{c}$ is the cosmic horizon. The specific forms of these four solutions are:
@@
$\tilde{\lambda}=$ $[972\Lambda M^{2}-72\Lambda^{3}AB-2\Lambda^{3}B^{3}+$
$\sqrt{(72\Lambda^{3}AB+2\Lambda^{3}B^{3}-972\Lambda M^{2})^{2}-4(\Lambda^{2}B^{2}-12\Lambda^{2}A)^{3}}]^{1/3},$ (101)
@@
$\tilde{\zeta}=$ $\sqrt{\frac{4B}{3}-\frac{(2)^{1/3}(\Lambda^{2}B^{2}-12\Lambda^{2}A)}{3\Lambda\tilde{\lambda}}-\frac{\tilde{\lambda}}{3(2)^{1/3}\Lambda}-\frac{12M}{\Lambda\tilde{\eta}}}.$ (103)
@@
$r_{i}=\frac{1}{2}\tilde{\eta}_{i}-\frac{1}{2}\tilde{\zeta}_{i}.$ (104)
Where:
$\tilde{\lambda}_{i}=$ $[972\Lambda M^{2}-72\Lambda^{3}AB-2\Lambda^{3}B^{3}+$
$\sqrt{(72\Lambda^{3}AB+2\Lambda^{3}B^{3}-972\Lambda M^{2})^{2}-4(\Lambda^{2}B^{2}-12\Lambda^{2}A)^{3}}]^{1/3},$ (105)
$\tilde{\eta}_{i}=$ $\sqrt{\frac{2B}{3}+\frac{(2)^{1/3}(\Lambda^{2}B^{2}-12\Lambda^{2}A)}{3\Lambda\tilde{\lambda}_{i}}+\frac{\tilde{\lambda}_{i}}{3(2)^{1/3}\Lambda}},$ (106)
$\tilde{\zeta}_{i}=$ $\sqrt{\frac{4B}{3}-\frac{(2)^{1/3}(\Lambda^{2}B^{2}-12\Lambda^{2}A)}{3\Lambda\tilde{\lambda}_{i}}-\frac{\tilde{\lambda}_{i}}{3(2)^{1/3}\Lambda}-\frac{12M}{\Lambda\tilde{\eta}_{i}}}.$ (107)
@@
$\tilde{\lambda}_{f}=$$\{972\Lambda(M-\omega)^{2}-72\Lambda^{3}A_{f}B_{f}-2\Lambda^{3}B_{f}^{3}+$
$\sqrt{[72\Lambda^{3}A_{f}B_{f}+2\Lambda^{3}B_{f}^{3}-972\Lambda(M-\omega)^{2}]^{2}-4(\Lambda^{2}B_{f}^{2}-12\Lambda^{2}A_{f})^{3}}\}^{1/3},$ (109)
@@
$\tilde{\zeta}_{f}=$ $\sqrt{\frac{4B_{f}}{3}-\frac{(2)^{1/3}(\Lambda^{2}B_{f}^{2}-12\Lambda^{2}A_{f})}{3\Lambda\tilde{\lambda}_{f}}-\frac{\tilde{\lambda}_{f}}{3(2)^{1/3}\Lambda}-\frac{12(M-\omega)}{\Lambda\tilde{\eta}_{f}}}.$ (111)
@@
$\tilde{\lambda}^{\prime}=$$\{972\Lambda(M-\omega^{\prime})^{2}-72\Lambda^{3}A^{\prime}B^{\prime}-2\Lambda^{3}B^{\prime 3}+$
$\sqrt{[72\Lambda^{3}A^{\prime}B^{\prime}+2\Lambda^{3}B^{\prime 3}-972\Lambda(M-\omega^{\prime})^{2}]^{2}-4(\Lambda^{2}B^{\prime 2}-12\Lambda^{2}A^{\prime})^{3}}\}^{1/3},$ (116)
@@
$\tilde{\zeta}^{\prime}=$ $\sqrt{\frac{4B^{\prime}}{3}-\frac{(2)^{1/3}(\Lambda^{2}B^{\prime 2}-12\Lambda^{2}A^{\prime})}{3\Lambda\tilde{\lambda}^{\prime}}-\frac{\tilde{\lambda}^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M-\omega^{\prime})}{\Lambda\tilde{\eta}^{\prime}}}.$ (118)
@@
$r_{i}^{\prime}=\frac{1}{2}\tilde{\eta}_{i}^{\prime}+\frac{1}{2}\tilde{\zeta}_{i}^{\prime}.$ (119)
Where:
$\tilde{\lambda}_{i}^{\prime}=$ $[972\Lambda M^{2}-72\Lambda^{3}AB-2\Lambda^{3}B^{3}+$
$\sqrt{(72\Lambda^{3}AB+2\Lambda^{3}B^{3}-972\Lambda M^{2})^{2}-4(\Lambda^{2}B^{2}-12\Lambda^{2}A)^{3}}]^{1/3},$ (120)
$\tilde{\eta}_{i}^{\prime}=$ $\sqrt{\frac{2B}{3}+\frac{(2)^{1/3}(\Lambda^{2}B^{2}-12\Lambda^{2}A)}{3\Lambda\tilde{\lambda}_{i}^{\prime}}+\frac{\tilde{\lambda}_{i}^{\prime}}{3(2)^{1/3}\Lambda}},$ (121)
$\tilde{\zeta}_{i}^{\prime}=$ $\sqrt{\frac{4B}{3}-\frac{(2)^{1/3}(\Lambda^{2}B^{2}-12\Lambda^{2}A)}{3\Lambda\tilde{\lambda}_{i}^{\prime}}-\frac{\tilde{\lambda}_{i}^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12M}{\Lambda\tilde{\eta}_{i}^{\prime}}}.$ (122)
@@
$\tilde{\lambda}_{f}^{\prime}=$$\{972\Lambda(M+\omega)^{2}-72\Lambda^{3}A_{f}^{\prime}B_{f}^{\prime}-2\Lambda^{3}B_{f}^{\prime 3}+$
$\sqrt{[72\Lambda^{3}A_{f}^{\prime}B_{f}^{\prime}+2\Lambda^{3}B_{f}^{\prime 3}-972\Lambda(M+\omega)^{2}]^{2}-4(\Lambda^{2}B_{f}^{\prime 2}-12\Lambda^{2}A_{f}^{\prime})^{3}}\}^{1/3},$ (124)
@@
$\tilde{\zeta}_{f}^{\prime}=$ $\sqrt{\frac{4B_{f}^{\prime}}{3}-\frac{(2)^{1/3}(\Lambda^{2}B_{f}^{\prime 2}-12\Lambda^{2}A_{f}^{\prime})}{3\Lambda\tilde{\lambda}_{f}^{\prime}}-\frac{\tilde{\lambda}_{f}^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M+\omega)}{\Lambda\tilde{\eta}_{f}^{\prime}}}.$ (126)
@@
$\tilde{\lambda}^{\prime\prime}=$$\{972\Lambda(M+\omega^{\prime})^{2}-72\Lambda^{3}A^{\prime\prime}B^{\prime\prime}-2\Lambda^{3}B^{\prime\prime 3}+$
$\sqrt{[72\Lambda^{3}A^{\prime\prime}B^{\prime\prime}+2\Lambda^{3}B^{\prime\prime 3}-972\Lambda(M+\omega^{\prime})^{2}]^{2}-4(\Lambda^{2}B^{\prime\prime 2}-12\Lambda^{2}A^{\prime\prime})^{3}}\}^{1/3},$ (131)
@@
$\tilde{\zeta}^{\prime\prime}=$ $\sqrt{\frac{4B^{\prime\prime}}{3}-\frac{(2)^{1/3}(\Lambda^{2}B^{\prime\prime 2}-12\Lambda^{2}A^{\prime\prime})}{3\Lambda\tilde{\lambda}^{\prime\prime}}-\frac{\tilde{\lambda}^{\prime\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M+\omega^{\prime})}{\Lambda\tilde{\eta}^{\prime\prime}}}.$ (133)
@@ 3.1 Painlevé coordinate and time-like geodesic line equation
Due to the fact that the constant time slice of the dragging coordinate system is not a flat Euclidean space in the radial direction, it is not the coordinate system we want. We adopt a method similar to ref. 24 to perform coordinate transformation again:
@@
Where $\hat{g}_{01}=\sqrt{\hat{g}_{00}(1-g_{11})}$. When a particle exits the event horizon of a black hole, (142) takes a positive sign. When a particle is incident from the cosmic horizon, (142) takes a negative sign. When considering self gravity, replace (142) with the following, $M\to M\mp\omega$, $Q_{h}\to Q_{h}-q_{h}$, $a=\frac{J}{M}\to\frac{J\mp j}{M\mp\omega}=\tilde{a}$. The minus sign and plus sign represent the outgoing particles and incoming particles, respectively.

### v2

@@ 2.1 Painlevé coordinate and time-like geodesic line equation
By substituting Eq. (46) and Eq. (47) into Eq. (13), we obtain the Painlevé coordinate:
@@
Obviously, in Eq. (48), the coordinate singularity at the event horizon is eliminated. The selection of positive and negative signs in Eq. (48) will have a clearer physical meaning in the equation of time-like geodesics.
In this article, we discuss particles that carry both electric and magnetic charges. So they are all massive particles, not following the optical geodesic equation. We consider massive particles as de Broglie s-waves, employing an approach analogous to that described in Ref. [28]. The time-like geodesic equation, derived from the WKB approximation, is presented below:
@@
Where $k$ is de Broglie wave number. By substituting $g_{00}$ and $g_{01}$ in Eq. (48) into Eq. (49), we can obtain:
@@
It can be clearly seen that when taking a positive sign, Eq. (52) is the equation of motion for the outgoing particles near the outer event horizon of the black hole, while when taking a negative sign, Eq. (52) is the equation of motion for the incoming particles near the cosmic horizon. Furthermore, if considering the influence of self gravity, $M$ in Eq. (48) and Eq. (52) should be replaced by $M\mp\omega$, and $Q_{h}$ in Eq. (48) and Eq. (52) should be replaced by $Q_{h}\mp q_{h}$. $\omega$ is the energy of the particle, $q_{h}$ is the equivalent charge of the particle. The minus sign and plus sign represent the outgoing particles and incoming particles, respectively.
@@ 2.2 Tunneling out of the black hole outer event horizon
According to Eq. (54), the imaginary part of the action can be written as:
@@
To proceed with our calculations, we substitute the Hamilton’s equation into Eq. (56):
@@
By substituting Eq. (57) and Eq. (58) into Eq. (56), we can obtain:
@@
Where $E_{Q_{h}}$ is the energy corresponding to the electromagnetic field. $M$, $E_{Q_{h}}$ are a fixed and unchanging quantity. $M-\omega$, $E_{Q_{h}-q_{h}}$ are the residual gravitational and electromagnetic energy of the black hole after radiation. It can be clearly seen that in our formula, energy conservation, charge conservation, and magnetic charge conservation are all reflected. It is worth emphasizing that in order to simplify the equivalent charge of a black hole after radiation to $Q_{h}-q_{h}$, we adopt the following assumption:
@@
Where $q_{e}$ and $q_{g}$ are the charges and magnetic charges of the emitted particles, respectively. As mentioned above, $Q_{h}$ is equivalent charge of black hole corresponding to density $\rho_{h}$. And $q_{h}$ is equivalent charge of the emitted particles also corresponding to density $\rho_{h}$. So we can get Eq. (60). The loss of gravitational energy and electromagnetic energy caused by particle tunneling in a black hole is represented by the following equations:
@@
By substituting Eq. (61) and Eq. (62) into Eq. (59), we can obtain:
@@
Substituting Eq. (64) into Eq. (63), we have:
@@ 2.3 Tunneling into the cosmic horizon
Eq. (72) and Eq. (73) respectively represent the changes in gravitational energy and electromagnetic energy after particles enter the cosmic horizon, from which it can be clearly seen that energy conservation, charge conservation, and magnetic charge conservation occur. Substituting Eq. (66), Eq. (72) and Eq. (73) into Eq. (71) yields:
@@ 2.4 Discussion
Comparing Eq. (80) and Eq. (81) with Eq. (66) and Eq. (75) respectively, the temperature $T^{\prime}$ at the event horizon surface of the black hole and the temperature $T^{\prime\prime}$ at the cosmic horizon surface are obtained:
@@ 3 Kerr-Newman-Kasuya de Sitter black hole
Let $A=\frac{3}{\Lambda}(a^{2}+Q_{h}^{2})$, $B=\frac{3}{\Lambda}-a^{2}$. The equation Eq. (90) has four solutions, $r_{non}$, $r_{-}$, $r_{+}$, $r_{c}$. Among them, $r_{non}$ is a negative value solution with no physical meaning, $r_{-}$ is the inner event horizon of black hole, $r_{+}$ is the outer event horizon of black hole, and $r_{c}$ is the cosmic horizon. The specific forms of these four solutions are:
@@
$\tilde{\zeta}=$ $\sqrt{\frac{4B}{3}-\frac{(2)^{1/3}(\Lambda^{2}B^{2}-12\Lambda^{2}A)}{3\Lambda\tilde{\lambda}}-\frac{\tilde{\lambda}}{3(2)^{1/3}\Lambda}-\frac{12M}{\Lambda\tilde{\eta}}},$ (96)
$\tilde{\lambda}=$ $\left[972\Lambda M^{2}-72\Lambda^{3}AB-2\Lambda^{3}B^{3}+\right.$
$\left.\sqrt{(72\Lambda^{3}AB+2\Lambda^{3}B^{3}-972\Lambda M^{2})^{2}-4(\Lambda^{2}B^{2}-12\Lambda^{2}A)^{3}}\right]^{1/3}.$ (97)
@@
$r_{i}=\frac{1}{2}\tilde{\eta}-\frac{1}{2}\tilde{\zeta}.$ (98)
@@
$\tilde{\zeta}_{f}=$ $\sqrt{\frac{4B_{f}}{3}-\frac{(2)^{1/3}(\Lambda^{2}B_{f}^{2}-12\Lambda^{2}A_{f})}{3\Lambda\tilde{\lambda}_{f}}-\frac{\tilde{\lambda}_{f}}{3(2)^{1/3}\Lambda}-\frac{12(M-\omega)}{\Lambda\tilde{\eta}_{f}}},$ (101)
$\tilde{\lambda}_{f}=$$\left\{972\Lambda(M-\omega)^{2}-72\Lambda^{3}A_{f}B_{f}-2\Lambda^{3}B_{f}^{3}+\right.$
$\left.\sqrt{[72\Lambda^{3}A_{f}B_{f}+2\Lambda^{3}B_{f}^{3}-972\Lambda(M-\omega)^{2}]^{2}-4(\Lambda^{2}B_{f}^{2}-12\Lambda^{2}A_{f})^{3}}\right\}^{1/3}.$ (102)
@@
$\tilde{\zeta}^{\prime}=$ $\sqrt{\frac{4B^{\prime}}{3}-\frac{(2)^{1/3}(\Lambda^{2}B^{\prime 2}-12\Lambda^{2}A^{\prime})}{3\Lambda\tilde{\lambda}^{\prime}}-\frac{\tilde{\lambda}^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M-\omega^{\prime})}{\Lambda\tilde{\eta}^{\prime}}},$ (108)
$\tilde{\lambda}^{\prime}=$$\left\{972\Lambda(M-\omega^{\prime})^{2}-72\Lambda^{3}A^{\prime}B^{\prime}-2\Lambda^{3}B^{\prime 3}+\right.$
$\left.\sqrt{[72\Lambda^{3}A^{\prime}B^{\prime}+2\Lambda^{3}B^{\prime 3}-972\Lambda(M-\omega^{\prime})^{2}]^{2}-4(\Lambda^{2}B^{\prime 2}-12\Lambda^{2}A^{\prime})^{3}}\right\}^{1/3}.$ (109)
@@
$r_{i}^{\prime}=\frac{1}{2}\tilde{\eta}+\frac{1}{2}\tilde{\zeta}.$ (110)
@@
$\tilde{\zeta}_{f}^{\prime}=$ $\sqrt{\frac{4B_{f}^{\prime}}{3}-\frac{(2)^{1/3}(\Lambda^{2}B_{f}^{\prime 2}-12\Lambda^{2}A_{f}^{\prime})}{3\Lambda\tilde{\lambda}_{f}^{\prime}}-\frac{\tilde{\lambda}_{f}^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M+\omega)}{\Lambda\tilde{\eta}_{f}^{\prime}}},$ (113)
$\tilde{\lambda}_{f}^{\prime}=$$\left\{972\Lambda(M+\omega)^{2}-72\Lambda^{3}A_{f}^{\prime}B_{f}^{\prime}-2\Lambda^{3}B_{f}^{\prime 3}+\right.$
$\left.\sqrt{[72\Lambda^{3}A_{f}^{\prime}B_{f}^{\prime}+2\Lambda^{3}B_{f}^{\prime 3}-972\Lambda(M+\omega)^{2}]^{2}-4(\Lambda^{2}B_{f}^{\prime 2}-12\Lambda^{2}A_{f}^{\prime})^{3}}\right\}^{1/3}.$ (114)
@@
$\tilde{\zeta}^{\prime\prime}=$ $\sqrt{\frac{4B^{\prime\prime}}{3}-\frac{(2)^{1/3}(\Lambda^{2}B^{\prime\prime 2}-12\Lambda^{2}A^{\prime\prime})}{3\Lambda\tilde{\lambda}^{\prime\prime}}-\frac{\tilde{\lambda}^{\prime\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M+\omega^{\prime})}{\Lambda\tilde{\eta}^{\prime\prime}}},$ (120)
$\tilde{\lambda}^{\prime\prime}=$$\left\{972\Lambda(M+\omega^{\prime})^{2}-72\Lambda^{3}A^{\prime\prime}B^{\prime\prime}-2\Lambda^{3}B^{\prime\prime 3}+\right.$
$\left.\sqrt{[72\Lambda^{3}A^{\prime\prime}B^{\prime\prime}+2\Lambda^{3}B^{\prime\prime 3}-972\Lambda(M+\omega^{\prime})^{2}]^{2}-4(\Lambda^{2}B^{\prime\prime 2}-12\Lambda^{2}A^{\prime\prime})^{3}}\right\}^{1/3}.$ (121)
@@ 3.1 Painlevé coordinate and time-like geodesic line equation
Due to the fact that the constant time slice of the dragging coordinate system is not a flat Euclidean space in the radial direction, it is not the coordinate system we want. We adopt a method similar to ref. [25] to perform coordinate transformation again:
@@
Where $\hat{g}_{01}=\sqrt{\hat{g}_{00}(1-g_{11})}$. When a particle exits the event horizon of a black hole, Eq. (130) takes a positive sign. When a particle is incident from the cosmic horizon, Eq. (130) takes a negative sign. When considering self gravity, replace Eq. (130) with the following, $M\to M\mp\omega$, $Q_{h}\to Q_{h}-q_{h}$, $a=\frac{J}{M}\to\frac{J\mp j}{M\mp\omega}=\tilde{a}$. The minus sign and plus sign represent the outgoing particles and incoming particles, respectively.